OpenStudy (anonymous):

SCIENTIFIC NOTATION: A. 234 B. .02 C. 1,234,000,000

OpenStudy (saifoo.khan):

2.34 x10^2 2 x10^-2 1.234x10^9

OpenStudy (saifoo.khan):

PLease hit >>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>>^^^^^^^^^^^^^^^^

OpenStudy (anonymous):

A. 234 = 2.34*10^2 ==== B. .02 = 2.0*10^-2 ==== C. 1,234,000,000 = 1.234 *109

OpenStudy (anonymous):

C. 1,234,000,000 = 1.234 *10^9

OpenStudy (saifoo.khan):

wow @ raheen, c part is great.

OpenStudy (saifoo.khan):

lol. now it;s fine.

OpenStudy (anonymous):

lool